A ship travels from New York to Southampton, England at 33 miles per hour. A plane traveling 605 miles per hour takes 104 hours less time to make the same trip. How far apart are the two cities?
Found 2 solutions by Paul, JWG:
Answer by Paul(988)

33(t+104)=605(t-104)
33t+3432=605t-62920
33t+3432-3432=605t-62920-3432
33t=605t-66352
33t-605t=605t-66352-605t
-572t=-66352 (divide both sides by -572)
x=116
Hence, the distance between the two cities is 116 miles

Answer by JWG(21)

Distance of ship=Distance of plane
Rate*Time=Distance
Ship: Rate=33mph, Time=x
Plane: Rate=605, Time=x-104
33x=605(x-104)
33x=605x-62920
572x=62920
x=110
Distance: 33*110=3,630
Double check: 605*(110-104)=605*6=3,630
3,630=3,630
Answer: 3,630